OpenStudy (anonymous):

f(x)=9/(4-radical(6x- -4). find the (implied) domain??

OpenStudy (across):

\[f(x)=\frac{9}{4-\sqrt{6x+4}}\]The denominator can't be zero, therefore: \[4-\sqrt{6x+4}=0,\]\[4=\sqrt{6x+4},\]\[16=6x+4,\]\[12=6x,\]\[x=2.\]The domain of x is then \[\left \{ x\in \mathbb{R}:x\neq 2 \right \}\]

jimthompson5910 (jim_thompson5910):

In addition, the radicand must be nonnegative So \[\large 6x+4\ge0\] \[\large 6x\ge-4\] \[\large x\ge-\frac{4}{6}\] \[\large x\ge-\frac{2}{3}\]